### v1

# Marine Saliency Segmenter: Object-Focused Conditional Diffusion with Region-Level Semantic Knowledge Distillation

Abstract
Marine Saliency Segmentation (MSS) plays a pivotal role in various vision-based marine exploration tasks. However, existing marine segmentation techniques face the dilemma of object mislocalization and imprecise boundaries due to the complex underwater environment. Meanwhile, despite the impressive performance of diffusion models in visual segmentation, there remains potential to further leverage contextual semantics to enhance feature learning of region-level salient objects, thereby improving segmentation outcomes. Building on this insight, we propose DiffMSS, a novel marine saliency segmenter based on the diffusion model, which utilizes semantic knowledge distillation to guide the segmentation of marine salient objects. Specifically, we design a region-word similarity matching mechanism to identify salient terms at the word level from the text descriptions. These high-level semantic features guide the conditional feature learning network in generating salient and accurate diffusion conditions with semantic knowledge distillation. To further refine the segmentation of fine-grained structures in unique marine organisms, we develop the dedicated consensus deterministic sampling to suppress overconfident missegmentations. Comprehensive experiments demonstrate the superior performance of DiffMSS over state-of-the-art methods in both quantitative and qualitative evaluations.
1 Introduction
Figure 1: Different from existing diffusion-based methods that directly condition on coarse-grained image-level visual or text features, our DiffMSS designs Region-Word Matching with Conditional Feature Learning Network (CFLN) and Semantic Knowledge Distillation (SKD) to capture fine-grained region-level visual features as accurate conditions for object-focused diffusion.
Marine Saliency Segmentation (MSS) focuses on segmenting visually salient objects within complex underwater environments to meet the growing requirement for fine-grained object recognition [12, 97, 38, 3]. Functionally, accurate recognition of marine instances contributes to applications like organism identification [10, 37], autonomous navigation [49, 100], and object detection [19, 4, 5]. However, the raw images captured directly by underwater vehicles tend to lose visual saliency, presenting various types of degradation, such as color distortion, low contrast, and blurred details [11, 22, 30, 96, 76]. Underwater degraded images with these defects usually exhibit indistinguishable object boundaries and a camouflaged appearance.
With advances in large-scale annotated datasets [18, 89] and deep network architectures, many saliency detection methods [83, 35, 64, 99, 70, 71, 87] in natural image domains have made remarkable performances. However, they face challenges in underwater environments, where the poor visibility and fine-grained structures of marine organisms (e.g., fish, corals) greatly degrade the accuracy [98]. Existing MSS methods [14, 25, 27, 20, 42] follow the basic paradigm of a learning-based backbone and a decoder for segmentation. However, they usually stack multiple convolutional sequences with limited representational power into deep networks to extract deep features that require a lot of computational resources [23, 19]. Without well-designed backbones, they remain vulnerable to visual degradation and suffer from inaccurate boundary segmentation.
Given the specific challenges posed by the MSS task, we explore the diffusion model [61] as a fitting solution due to its strong generative capabilities with well-defined conditional inputs. Despite the impressive performance of diffusion models with conditional prompts in visual segmentation tasks [59, 55, 7, 77], the potential of leveraging contextual semantics to generate diffusion conditions remains underexplored. While the contextual semantics usually come from the caption of the image, word-level concepts within the text can provide more direct fine-grained information about the salient object [74, 95]. In fact, by extracting and aligning these word-level semantic and visual features, we can guide the diffusion model to focus on these key salient regions and improve its segmentation accuracy.
Motivated by the aforementioned analysis, we propose DiffMSS, an innovative diffusion-based marine saliency segmenter that leverages a region-level knowledge distillation scheme to guide the segmentation of marine salient objects. As depicted in Fig. 2, we first design a Word-level Semantic Saliency Extraction (WSSE) to adaptively identify salient terms described in the given text through region-word similarity matching. Then, these high-level text features transfer contextual semantic information to the Conditional Feature Learning Network (CFLN) based on Semantic Knowledge Distillation (SKD), guiding it to generate region-level visual features as object-focused diffusion conditions. This distillation scheme enables DiffMSS to improve inference efficiency without relying on WSSE and SKD for semantic saliency transfer during the testing phase. To further refine the segmentation of fine-grained structures, we develop a dedicated Consensus Deterministic Sampling (CDS) to suppress inaccurate segmentation caused by overconfidence in camouflaged marine instances.
Our key contributions are summarized as follows:
  • We propose DiffMSS, a novel object-focused diffusion model for marine saliency segmentation. It simplifies the challenging MSS task into a series of identification, segmentation, and refinement procedures.
  • We design region-level semantic knowledge distillation to capture fine-grained visual features as guiding conditions for object-focused diffusion. We also propose a dedicated CDS scheme to suppress overconfident missegmentations in camouflaged instances.
  • Comprehensive experiments on the public datasets validate that our DiffMSS surpasses existing state-of-the-art solutions in both qualitative and quantitative outcomes.
Figure 2: Overview of the proposed DiffMSS. Its training mainly contains three procedures: (a) Given the image caption, Word-level Semantic Saliency Extraction identifies salient concepts in terms of words via region-word matching. (b) Semantic Knowledge Distillation transfers the identified word-level semantic tokens into the Conditional Feature Learning Network to generate region-level conditions for object-focused diffusion. (c) Consensus Sampling enables fine-grained structural segmentation of intricate marine instances via deterministic ensemble scheme. Note the top green box (i.e., Word-level Semantic Saliency Extraction) and the Semantic Knowledge Distillation in the bottom yellow box are utilized exclusively during training and will deactivated in the testing phase.
2 Related Work
2.1 Marine Saliency Segmenatation
Existing MSS methods can be roughly divided into handcrafted feature-based methods [28, 56, 19, 29, 86, 32] and deep learning-based methods [31, 27, 25, 23, 79, 80, 38]. Early handcrafted feature-based methods relied on low-level visual features to achieve segmentation [9, 53]. With the rise and advancement of visual foundation models, various network architectures have been proposed to address MSS. Li et al. [38] proposed a feature interaction encoder and cascaded decoder to extract more comprehensive information, while Liu et al. [44] combined channel and spatial attention modules to refine feature maps to obtain better object boundaries. Although these CNN-based models are effective, they cannot capture the long-range dependencies of complex marine objects and ignore the connectivity between discrete pixels [45]. Recently, instead of linearly stacking multiple convolutional layers in the network, several deep learning-based USD methods [2, 23, 43, 21, 14] incorporated visual transformers with wider receptive fields into their deep architectures. This way alleviates the computational burden brought by convolution to some extent, but these methods are unreliable in capturing saliency information by improving the encoder architecture.
2.2 Text-supervised Feature Matching
With the rise of text-supervised semantic segmentation, many studies [82, 74, 40, 78, 57, 81, 1] have utilized text prompts to enhance segmentation performance. These large vision-language models [58, 39, 41, 63] have been trained on large text-image datasets such as LAION-5B [60], enabling them to understand the alignment between text descriptions and visual elements. They train image and text encoders to align image-text pairs within a joint embedding space, thereby generating segmentation results with zero-shot supervision [52, 13, 94, 93]. Although straightforward, images may contain multiple object instances, and the semantic features of the text should match corresponding segments rather than the entire image. Several region-text alignment methods [34, 62, 1] have been proposed to strengthen the consistency between the segmented region and the text description, which enables the network to focus on segmenting the relevant regions described in the text.
2.3 Diffusion-based Image Segmentation
With their powerful generative capabilities, diffusion models have achieved impressive performance in terms of image restoration [90, 91, 65], object detection [6, 72, 92, 73], and depth estimation [84, 51, 33]. By leveraging the adaptive characteristics of the diffusion process, diffusion models have shown potential in various segmentation tasks [66, 24, 75, 47, 55]. For instance, DiffuMask [75] utilizes cross-modal attention maps between image features and conditional text embeddings to segment the most prominent object indicated by text prompts. LD-ZNet [55] performs text-based synthetic image segmentation by revealing rich semantic information within its internal features. However, existing diffusion models employ pixel-level corruption to generate the noised mask directly from the GT, which causes the model to mistakenly assume that the restored contours from the noised mask are accurate [67]. In addition, these methods often produce conditional features with limited discriminative representation. To address this, we propose a conditional feature learning network under the guidance of region-level semantic knowledge distillation to robustly generate discriminative conditional features.
3 Methodology
We first introduce the word-level semantic saliency extraction for identifying words that describe salient objects in Section 3.1. Then, Section 3.2 presents the semantic knowledge distillation for guiding the conditional feature learning network to generate region-level features as conditions in diffusion. Finally, we describe object-focused conditional diffusion and consensus deterministic sampling for segmenting fine-grained masks in Section 3.3.
3.1 Word-Level Semantic Saliency Extraction via Word-Region Matching
Unlike image-text alignment, region-word matching focuses on aligning segmented regions (rather than the whole image) with words in a joint embedding space. It ensures consistency between the segmented region and textual description by learning key salient objects in the image.
Image-Text Segmenter. We first introduce an image segmenter and a text segmenter: the former decomposes an image into region segments, while the latter decomposes a text into word segments. It enables both the image and text segmenters to learn region-word consensus when segmenting the input image $\mathcal{I}$ with a paired text $\mathcal{T}$. Specifically, given an image $\mathcal{I}\in\mathbb{R}^{H\times W\times C_{v}}$ and the corresponding text $\mathcal{T}\in\mathbb{R}^{N_{t}\times C_{t}}$, where $H,W,C_{v}$ represent the height, width, channel of image $\mathcal{I}$, and $N_{t},C_{t}$ represent the number, dimension of the words. We utilize the image segmenter and text segmenter to process the image-text pairs, thus obtaining a group of $M$ region masks $\mathbb{X}^{v}=\left\{\mathcal{X}_{i}^{v}\right\}_{i=1}^{M}$ and the corresponding text $\mathbb{X}^{t}=\left\{\mathcal{X}_{j}^{t}\right\}_{j=1}^{N}$ of $N$ single-word nouns. That is, $\mathbb{X}^{v}$ contains several sub-images $\mathcal{X}_{i}^{v}$ obtained by cropping and masking relevant regions from the input image $\mathcal{I}$, while $\mathbb{X}^{t}$ takes a text $\mathcal{T}$ of length $N$ as input and extracts each word $\mathcal{X}_{j}^{t}$ in $\mathcal{T}$.
Figure 3: Illustration of the Conditional Features Learning Network (CFLN). It cascades the input image $\mathcal{I}$, intermediate sample $\mathbf{x}_{t}$, and time token $t$ through four Transformer-based blocks to extract region-level features as conditions in diffusion.
Saliency Word-Token Discover. We then select high-confidence salient words as the guided tokens, instead of the whole caption. We employ the image encoder $E_{v}(\mathcal{X}_{i}^{v})$ and text encoder $E_{t}(\mathcal{X}_{j}^{t})$ of the pre-trained CLIP model [58] to extract semantic saliency features from the highlighted regions and words, respectively. For each input image, the visual embedding tokens $\mathcal{F}_{i}^{v}$ are calculated as follows:
$\mathcal{F}_{i}^{v}=\mathcal{W}^{v}\times\mathcal{Z}^{v}_{i},i\in\left\{1,2,\dots,M\right\},$ (1)
where $\mathcal{Z}^{v}_{i}$ represents the visual features provided by the image encoder $\mathcal{Z}^{v}_{i}=E_{v}(\mathcal{X}_{i}^{v})$, and $\mathcal{W}^{v}$ is the projection matrix that converts $\mathcal{Z}^{v}_{i}$ into the vision embedding tokens $\mathcal{F}_{i}^{v}$. In the same way, the word prompts $\mathcal{X}_{j}^{t}$ are transformed into textual embedding tokens $\mathcal{F}_{j}^{t}$ through the projection matrix $\mathcal{W}^{t}$. Both types of tokens have the same dimensionality in the joint embedding space and the region-word similarity is calculated as follows:
$R_{k}=\frac{1}{MN}\sum_{i=1}^{M}\sum_{j=1}^{N}Softmax\left(\mathcal{F}_{i}^{v}{\mathcal{F}_{j}^{t}}^{T}\right).$ (2)
After that, we calculate the average $m=\text{Mean}(R_{k})$ of obtained similarity scores and then select the indices of scorers exceeding $m$ from the candidate list containing $L_{t}$ tokens with priority. Mathematically, it is expressed as:
$\mathcal{F}^{st}=\left\{k\mid R\{k\}\geq m,\,k\in\left\{1,2,\dots,L_{t}\right\}\right\},$ (3)
where $\mathcal{F}^{st}$ defines the index of selected tokens from the candidate list with the score priority. To avoid region-word mismatches, we use the noun-selector [15] to filter the segmented words. That is, some words that are irrelevant to the salient objects in the visual domain, such as prepositions and pronouns, are not considered for guiding the conditional feature generation.
3.2 Region-Level Semantic Saliency Knowledge Distillation
By leveraging high-level semantic tokens and aligning them with visual features, it can guide the diffusion model to focus on the salient regions, thereby improving the accuracy of object segmentation.
Conditional Features Learning Network. The network aims to generate region-level conditions that enable the diffusion model to effectively identify salient objects at each denoising step. Its design needs to meet three requirements: 1) Extracting discriminative salient features based on the image content; 2) Providing region-level conditions associated with the current denoising step; and 3) Capturing long-range dependencies and contextual information of the whole image. In addition, the inherent degradation characteristics of underwater images significantly interfere with the extraction of discriminative image features, thus diminishing the performance of the mask decoder.
To address this issue, we design a well-generalized Conditional Feature Learning Network (CFLN) based on the pyramid vision transformer [69]. As shown in Fig. 3, it extracts visual features $\{\mathcal{F}_{l}\}_{l=1}^{4}$ from a triplet data $(\mathcal{I},\mathbf{x}_{t},t)$, in which $\mathcal{I}$ represents the input image, $\mathbf{x}_{t}$ denotes previous sampling results, and $t$ represents the denoising step. $\mathbf{x}_{t}$ serves as a guiding cue to assist CFLN in adaptively focusing on specific regions, while adding the time step $t$ aims to improve the synchronization of the extracted conditions in the denoising step. To achieve this, we employ the zero overlap embedding to incorporate the noise mask $\mathbf{x}_{t}$ into the first block in a controlled manner without disrupting the Transformer structure, which is expressed as follows:
$\mathcal{F}_{l}=\begin{cases}Norm\left(\mathbb{R}\left({Conv}(\mathcal{I})+{Conv}_{z}(\mathbf{x}_{t})\right)\right),&l=1,\\ Norm\left(\mathbb{R}({Conv}(\mathcal{F}_{l-1}))\right),&l\neq 1.\end{cases}$ (4)
where ${Conv}(\mathcal{I})$ and ${Conv}_{z}(\mathbf{x}_{t})$ denote convolutional layers, differing in whether the weights and biases are initialized to zero. $Norm(\cdot)$ denotes layer normalization, while $\mathbb{R}(\cdot)$ represents transforming the feature map into tokens.
In addition to embedding the noise mask, we desire that the CFLN can adaptively tune the conditional features over time steps. We propose a scheme to concatenate the time token $t$ with the embedding patches $\mathcal{F}_{l}$, as follows:
$\mathcal{F}_{l}^{v}=\mathbb{R}^{-1}\left(MHA([t;\mathcal{F}_{l}])\right),l\in\left\{1,2,3,4\right\},$ (5)
where [;][;][ ; ] refers to the connection operation, $\mathbb{R}^{-1}$ reconverts tokens into multi-scale features, and $MHA$ represents the multi-head attention.
Word-level Knowledge Transfer. In Section 3.1, we have obtained word-level tokens that contain semantic information, which assists in identifying salient objects within the entire image. Based on this, we utilize the Semantic Knowledge Distillation (SKD) to constrain the generation of diffusive conditions throughout the training phase. Specifically, we design two distinct projectors to map the textual features of tokens (denoted as $\mathcal{F}^{st}$) and the visual features of conditions (denoted as $\mathcal{F}_{l}^{v}$) into a unified latent feature space. In other words, $\mathcal{F}^{st}$ are selected word-level tokens derived from the text encoder, while $\mathcal{F}_{l}^{v}$ represents the conditional features generated by the CFLN module. Considering that these two features should exhibit consistency across the latent space, we define a consistent loss $\mathcal{L}_{\textit{consist}}$ to constrain them, expressed as:
$\mathcal{L}_{\textit{consist}}=1-\frac{{Proj}(\mathcal{F}_{l}^{v})_{v}\cdot{Proj}(\mathcal{F}^{st})_{t}}{\|{Proj}(\mathcal{F}_{l}^{v})_{v}\|\,\|{Proj}(\mathcal{F}^{st})_{t}\|},$ (6)
where ${Proj}(\cdot)_{t}$ and ${Proj}(\cdot)_{v}$ represent the projectors for mapping textual tokens and conditional features into the latent embedding space, respectively.
For the discrete features, we employ the Local Emphasis (LE) module in [68] and convolutional calculation to aggregate them, expressed as follows:
$\mathcal{F}_{l}^{v}=Conv([\mathcal{F}_{l+1}^{v},LE(\mathcal{F}_{l}^{v})]),l\in\{3,2,1\},$ (7)
where the aggregated feature is defined as $\mathcal{F}_{a}^{v}=LE(\mathcal{F}_{4}^{v})$ and serves as the region-level diffusion conditions.
3.3 Object-Focused Diffusion and Sampling
Compared to traditional segmentation baselines, our proposed DiffMSS framework employs a revised conditional diffusion model with feature consistency learning to generate predicted masks. However, iterative diffusion and sampling may face two inherent challenges when generating masks: 1) Restoring a high-fidelity mask from low signal-to-noise ratio noise based on visual features is challenging; 2) Degraded images may cause well-trained models to produce occasional missegmentations due to overconfidence. The reason for this dilemma is that the model tends to choose the path of least resistance for parameter learning. They rely on more obvious noise masks instead of utilizing conditional features for generation. To address these issues, we propose the Object-Focused Conditional Diffusion (OFCD) and Consensus Deterministic Sampling (CDS) to achieve fine-grained structural segmentation for marine camouflage objects.
Object-Focused Conditional Diffusion. In forward diffusion, given a training sample $x_{0}\sim q(x_{0})$, the noised samples $\{x_{t}\}_{t=1}^{T}$ are obtained according to the following Markov process:
$q(\mathrm{x}_{t}|\mathrm{x}_{t-1})=\mathcal{N}\left(\mathrm{x}_{t};\sqrt{1-\beta_{t}}\mathrm{x}_{t-1},\beta_{t}\mathbf{I}\right),$ (8)
where $\beta_{t}$ denote the pre-defined noise schedule at $t$-th time step. The marginal distribution of $x_{t}$ can be described as:
$q(\mathrm{x}_{1:T}|\mathrm{x}_{0})=\mathcal{N}\left(\mathrm{x}_{t};\sqrt{\bar{\alpha}_{t}}\mathrm{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I}\right),$ (9)
where $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}={\textstyle\prod_{i=1}^{t}\alpha_{i}}$.
Figure 4: The conditional feature maps and mask predictions at different sampling steps $t$.
The previous diffusion paradigm that learns a conditional reverse process $p_{\theta}(\mathrm{x}_{T:0}|\mathrm{y})$ without modifying the forward diffusion $q(\mathrm{x}_{1:T}|\mathrm{x}_{0})$, ensuring the sampled $\hat{\mathrm{x}}_{0}$ is faithful to the raw data distribution. Instead of taking input image $\mathrm{y}$ as an invariant condition, we employ the aggregated features $\mathcal{F}_{a}^{v}$ produced by the CFLN module as conditions. Mathematically, it is expressed as follows:
$q(\mathrm{x_{t-1}|x}_{t},\mathcal{F}_{a}^{v})=\mathcal{N}\left(\mathrm{x}_{t-1};\mu_{\theta}(\mathrm{x}_{t},\mathcal{F}_{a}^{v},t),\delta^{2}_{t}\mathbf{I}\right),$ (10)
where the variance $\delta^{2}_{t}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}$, and the mean $\mu_{\theta}$ is defined as follows:
$\mu_{\theta}(\mathrm{x}_{t},\mathcal{F}_{a}^{v},t)=\frac{1}{\sqrt{\bar{\alpha}_{t}}}(\mathrm{x}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(\mathrm{x}_{t},\mathcal{F}_{a}^{v},t)),$ (11)
where $\epsilon_{\theta}(\mathrm{x}_{t},\mathcal{F}_{a}^{v},t)$ represents the predicted noise by optimizing the parameters $\theta$ of our proposed DiffMSS model. We transform the estimated noise $\epsilon_{\theta}$ into the salient mask conditioned on the region-level aggregated features $\mathcal{F}_{a}^{v}$, which is defined as follows:
$\hat{\mathrm{x}}_{0}=\frac{\mathrm{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\,\epsilon_{\theta}(\mathrm{x}_{t},\mathcal{F}_{a}^{v},t)}{\sqrt{\bar{\alpha}_{t}}},$ (12)
where $\hat{\mathrm{x}}_{0}$ is predicted by our model $f_{\theta}(\mathrm{x}_{t},\mathcal{F}_{a}^{v},t)$. Based on this, we utilize the saliency mask $\mathrm{x}_{0}$ corresponding to the real-world underwater scene as a reference to constrain the rationality of the predicted mask, as expressed below:
$\mathcal{L}_{\textit{mask}}=\mathcal{L}_{\textit{BCE}}^{w}(\hat{\mathrm{x}}_{0},\mathrm{x}_{0})+\mathcal{L}_{\textit{IoU}}^{w}(\hat{\mathrm{x}}_{0},\mathrm{x}_{0}).$ (13)
Based on the semantic knowledge distillation term $\mathcal{L}_{consist}$ and saliency mask refinement term $\mathcal{L}_{mask}$, the hybrid objective function $\mathcal{L}_{total}$ is defined by combining them as follows:
$\mathcal{L}_{Total}=\mathcal{L}_{consist}+\lambda\mathcal{L}_{mask},$ (14)
where $\lambda=0.5$ is weighted to coordinate the significance of each term in the experiment.
Figure 5: Visualization comparisons between our DiffMSS and state-of-the-art methods on the common underwater salient objects. The segmentation results are marked in orange.
Figure 6: Visualization comparisons between our DiffMSS and state-of-the-art methods on the challenging underwater camouflage objects with fine-grained structures. The segmentation results are marked in orange.
To illustrate that DiffMSS can reduce noise and progressively focus on salient objects, we display the predicted results and conditional feature maps captured at different sampling steps in Fig. 4. It is clear that the model progressively focuses on salient objects and refines the mask, enabling it to establish well-defined boundaries based on the foreground objects.
Consensus Deterministic Sampling. To improve the segmentation accuracy of fine-grained anatomical structures in marine instances, we introduce a Consensus Deterministic Sampling (CDS) method to aggregate predictions from each denoising step, which is inspired by the saliency detection annotation in [85]. Specifically, we denote the denoised image $\hat{\mathrm{x}}_{0}$ as $P_{t}$ at each sampling stage $t$. After obtaining multiple predictions $\{P_{t}\}_{t=1}^{T}$, which are then calculated as binary masks by setting an average threshold. These predictions $\{P_{t}^{b}\}_{t=1}^{T}$ vote on the position of each point to generate a candidate mask. The probability value of each selected point is calculated as the average of all predictions. Mathematically, it is defined as follows:
$M_{pre}=\left\lfloor\frac{1}{T}\sum_{t=1}^{T}P_{t}^{b}+\varphi\right\rfloor\ast Norm\left(\frac{1}{T}\sum_{t=1}^{T}P_{t}\right),$ (15)
where $\varphi=0.5$ represents the average threshold calculated with samples. The CDS schedule generates multiple predictions by iterative sampling, which enables us to improve mask accuracy through ensemble techniques.
Training and Inference. The training phase of DiffMSS requires the degraded image $\mathcal{I}$, the corresponding caption $\mathcal{T}$, and the reference saliency mask $\mathrm{x}_{0}$ for supervision, whereas its inference only requires the degraded image $\mathcal{I}$ as input. In other words, the inference of DiffMSS relies solely on Object-Focused Conditional Diffusion (OFCD) to generate segmentation results, without the Word-level Semantic Saliency Extraction (WSSE) procedure. More detailed training and sampling procedures can be found in the supplementary material.
4 Experiments
4.1 Experimental Setups
Implementation Details. The proposed DiffMSS is trained using the Pytorch framework on two NVIDIA GeForce RTX 4090 GPUs for $150$ epochs. During the training phase, the batch size and patch size are set to 32 and $256\times 256$, respectively. The Adam optimizer comes with an initial learning rate of $1\times 10^{-4}$ and decreases it by a factor of 0.8 after every ten epochs. The diffusion steps are set to $T=1000$ with a noise schedule $\beta_{t}$ that increases linearly from 0.0001 to 0.02, while the sampling steps are set to $S=10$ for efficient restoration. More detailed hyperparameter settings can be found in the supplementary material.
Config. USOD10K SUIM UFO-120
$F_{\beta}^{w}\uparrow$ $E_{\phi}^{m}\uparrow$ $S_{\alpha}\uparrow$ $m_{IoU}\uparrow$ $M_{AE}$downarrow $F_{\beta}^{w}\uparrow$ $E_{\phi}^{m}\uparrow$ $S_{\alpha}\uparrow$ $m_{IoU}\uparrow$ $M_{AE}$$\downarrow$ $F_{\beta}^{w}\uparrow$ $E_{\phi}^{m}\uparrow$ $S_{\alpha}\uparrow$ $m_{IoU}\uparrow$ $M_{AE}$downarrow
SUIM-Net [25] 0.783 0.856 0.797 0.628 0.1011 0.807 0.867 0.826 0.705 0.0787 0.711 0.733 0.679 0.639 0.1326
RMFormer [14] 0.828 0.910 0.867 0.746 0.0439 0.830 0.908 0.859 0.753 0.0623 0.790 0.851 0.796 0.711 0.1085
BBRF [48] 0.799 0.875 0.853 0.731 0.0462 0.816 0.891 0.856 0.748 0.0679 0.739 0.746 0.759 0.683 0.1119
TC-USOD [23] 0.907 0.968 0.922 0.872 0.0201 0.879 0.951 0.893 0.827 0.0388 0.856 0.917 0.859 0.816 0.0631
RobustSAM [8] 0.872 0.947 0.908 0.861 0.0396 0.859 0.918 0.865 0.806 0.0573 0.839 0.893 0.847 0.802 0.0717
DualSAM [88] 0.917 0.968 0.924 0.881 0.0185 0.876 0.937 0.881 0.817 0.0465 0.858 0.921 0.861 0.813 0.0637
DiffMSS (Ours) 0.923 0.967 0.932 0.893 0.0189 0.891 0.947 0.908 0.835 0.0376 0.867 0.927 0.873 0.821 0.0566
Table 1: Quantitative evaluation of our DiffMSS and state-of-the-art methods on three public underwater datasets (USOD10K [23], SUIM [25], and UFO-120 [26]). The best and second-best results are highlighted with bold and underlined, respectively.
Benchmark Datasets. We evaluate DiffMSS on three popular USOD benchmarks (USOD10K [23], SUIM [25] and UFO-120 [26]), all of which are real-world underwater images with references. Specifically, we follow the default settings in USOD10K, using 7,178 images for training and 1,026 images for testing. Meanwhile, we use 1,300 images from each of the SUIM and UFO-120 datasets for training, with the remaining images reserved for testing, respectively. For a fair comparison, all compared methods are retrained on the same data with their default settings.
Evaluation Metrics. We adopt five commonly used metrics for MSS tasks evaluation, including weighted F-measure ($F_{\beta}^{w}$) [50], max E-measure ($E_{\phi}^{m}$) [17], S-measure ($S_{\alpha}$) [16], Mean Intersection over Union ($m_{IoU}$) [46], and mean absolute error ($M_{AE}$) [54].
4.2 Comparison with State-of-the-Arts
[…]
SUIM-Net [25] 12.22 71.46 0.265 0.1035
RMFormer [14] 174.19 563.14 0.315 0.0598
BBRF [48] 74.01 31.13 0.107 0.0629
TC-USOD [23] 117.64 29.64 0.089 0.0318
RobustSAM [8] 407.76 1492.60 0.137 0.0490
DualSAM [88] 159.95 318.05 0.081 0.0322
DiffMSS (Ours) 68.41 24.98 0.033 0.0296
[…]
We compare the proposed DiffMSS model with six state-of-the-art (SOTA) saliency object detection methods, including SUIM-Net [25], RMFormer [14], BBRF [48], TC-USOD [23], RobustSAM [8], and DualSAM [88].
Qualitative Evaluation. As shown in Fig. 5, we first conduct a visual comparison of our DiffMSS with several SOTA methods on two common underwater salient objects. Compared with SUIM-Net, RobustSAM, and DualSAM, our model shows superior segmentation performance, especially involving objects with blurred boundaries. We then evaluate DiffMSS on the challenging marine camouflage objects with fine-grained structures. As shown in Fig. 6, our model consistently achieves superior segmentation results, characterized by well-defined boundaries and strong robustness against underwater noise and artifacts.
Quantitative Evaluation. We further conduct a quantitative evaluation of these compared methods, and the results are presented in Table 1. DiffMSS consistently achieves the best or second-best scores across all metrics on the three public underwater datasets, especially performing well on UFO-120. This demonstrates the robustness and generalization ability of our model in handling marine saliency segmentation tasks under various challenging conditions.
4.3 Evaluation of Model Efficiency
Parameters and FLOPs. Considering the limited computational resources of underwater embedded devices, our DiffMSS ensures segmentation accuracy while excelling in terms of parameters and FLOPs. As shown in Table 2, DiffMSS’s 68.41M parameters are lower than RMFormer (174.19M) and RobustSAM (407.76M). Moreover, our DiffMSS achieves the lowest FLOPs (24.98G) that outperforms other methods like SUIM-Net and TC-USOD, making it a more efficient choice for underwater applications with limited computing resources.
Figure 7: Visual ablation of segmentation results with (-w/ SKD) and without (-w/o SKD) Semantic Knowledge Distillation.
Figure 8: Conditional features ablations of with (-w/ SKD) and without (-w/o SKD) Semantic Knowledge Distillation at different sampling steps.
-w/o SKD -w/ SKD $F_{\beta}^{w}\uparrow$ $E_{\phi}^{m}\uparrow$ $S_{\alpha}\uparrow$ $M_{AE}$downarrow
$\checkmark$ 0.860 0.932 0.905 0.1110
$\checkmark$ 0.923 0.967 0.932 0.0189
Table 3: Ablation study of Semantic Knowledge Distillation.
Inference Time. Unlike these compared methods that stack multiple convolutional sequences or Segment Anything Model (SAM)-based, our DiffMSS exploits object-focused conditional diffusion to optimize computational efficiency while maintaining effective deep feature extraction. As shown in Table 2, DiffMSS achieves the fastest inference time of 0.033s. Although DualSAM’s inference time is relatively short (0.081s), it is still more than twice ours. The efficiency is mainly attributed to the semantic knowledge distillation that transfers high-level text semantic information, and the inference requires ten sampling steps to generate predicted results from a single input image.
I-T Match. R-W Match. $F_{\beta}^{w}\uparrow$ $E_{\phi}^{m}\uparrow$ $S_{\alpha}\uparrow$ $M_{AE}$downarrow
$\checkmark$ 0.887 0.943 0.917 0.0789
$\checkmark$ 0.923 0.967 0.932 0.0189
Table 4: Ablation study of different modality matching schemes for semantic knowledge distillation. “I-T” represents image-text matching, while “R-W” represents region-word matching.
4.4 Ablation Study
Ablation Study of Semantic Knowledge Distillation. We first conduct a visual ablation study with and without semantic knowledge distillation (denoted as “-w/ SKD” and “-w/o SKD”). Fig. 7 shows that the absence of caption guidance in “-w/o SKD” leads to over-segmentation or under-segmentation, while Fig. 8 presents that “-w/ SKD” contributes to generating fine-grained conditional features. In the quantitative evaluation in Table 3, “-w SKD” outperforms “-w/o SKD” in four indicators, which indicates the necessity of SKD for salient object segmentation.
Ablation Study of Different Modality Matching. In the semantic saliency extraction procedure, we further conduct an ablation study on different modality matching in the “-w/ SKD” case, including Image-Text matching (denoted as “I-T”) or Region-Word matching (denoted as “R-W”). As shown in Table 4, compared with “-w/o SKD”, “I-T” matching can improve the performance of saliency segmentation to a certain extent, but our “R-W” matching scheme achieves the highest scores across all metrics, which demonstrates the effectiveness of word-level semantic alignment in achieving object-focused diffusion.
Necessity of Features Aggregation in CFLN. Table 5 presents a discussion on the impact of aggregating different layer features ($\mathcal{F}_{1}$, $\mathcal{F}_{2}$, $\mathcal{F}_{3}$, $\mathcal{F}_{4}$) as diffusion conditions in the CFLN module. The scores of the four metrics gradually increase with the aggregation of more feature layers. All features are aggregated together to produce the highest $F_{\beta}^{w}$, $E_{\phi}^{m}$, and $S_{\alpha}$, along with the lowest $M_{AE}$, while the number of model parameters and computational burden increase only slightly compared to the former.
$\mathcal{F}_{1}$ $\mathcal{F}_{2}$ $\mathcal{F}_{3}$ $\mathcal{F}_{4}$ $F_{\beta}^{w}\uparrow$ $E_{\phi}^{m}\uparrow$ $S_{\alpha}\uparrow$ $M_{AE}$downarrow Param. FLOPs
$\checkmark$ 0.664 0.853 0.796 0.0719 65.23 19.49
$\checkmark$ $\checkmark$ 0.831 0.943 0.899 0.0399 66.62 20.31
$\checkmark$ $\checkmark$ $\checkmark$ 0.883 0.955 0.908 0.0229 67.59 21.66
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 0.923 0.967 0.932 0.0189 68.41 24.98
Table 5: Ablation study of aggregating different layer features as region-level diffusion conditions in CFLN module.
Complementarity of Loss Function. Table 6 presents a discussion on various loss functions, including $\mathcal{L}_{\text{consist}}$, $\mathcal{L}_{\text{BCE}}^{w}$, and $\mathcal{L}_{\text{IoU}}^{w}$. When using only $\mathcal{L}_{\text{consist}}$, the model produced the lowest scores, with $F_{\beta}^{w}$ at 0.514 and $M_{AE}$ at 0.2489. While semantic knowledge effectively supports saliency localization, it remains insufficient for precise segmentation of object boundaries. Adding $\mathcal{L}_{\text{BCE}}^{w}$ significantly enhanced performance, raising $F_{\beta}^{w}$ to 0.856 and lowering $M_{AE}$ to 0.1369. The model achieved optimal scores when all three loss functions were combined, suggesting that these loss functions complement one another to deliver the most efficient model performance across all metrics.
$\mathcal{L}_{consist}$ $\mathcal{L}_{BCE}^{w}$ $\mathcal{L}_{IoU}^{w}$ $F_{\beta}^{w}\uparrow$ $E_{\phi}^{m}\uparrow$ $S_{\alpha}\uparrow$ $M_{AE}\downarrow$
$\checkmark$ 0.514 0.775 0.656 0.2489
$\checkmark$ $\checkmark$ 0.860 0.932 0.895 0.1109
$\checkmark$ $\checkmark$ 0.892 0.951 0.919 0.0247
$\checkmark$ $\checkmark$ 0.856 0.925 0.873 0.1369
$\checkmark$ $\checkmark$ $\checkmark$ 0.923 0.967 0.932 0.0189
Table 6: Ablation study of loss function terms.
Effectiveness of Consensus Deterministic Sampling. Table 7 presents a discussion on the impact of CDS scheme for saliency segmentation. Without CDS scheme (“-w/o CDS”), the model achieves lower scores on all four evaluation metrics. In contrast, with CDS scheme (“-w/ CDS”), the scores improved across all metrics, indicating a significant enhancement in segmentation accuracy and a reduction in errors. We further perform a visual comparison between the two cases. As shown in Figure 9, it can be seen that the CDS scheme significantly improves the fine-grained segmentation performance of marine instances.
-w/o CDS -w/ CDS $F_{\beta}^{w}\uparrow$ $E_{\phi}^{m}\uparrow$ $S_{\alpha}\uparrow$ $M_{AE}$downarrow
$\checkmark$ 0.886 0.935 0.906 0.0274
$\checkmark$ 0.923 0.967 0.932 0.0189
Table 7: Ablation study of Consensus Deterministic Sampling.
Figure 9: Visual ablation of Consensus Deterministic Sampling.
5 Conclusion
In this paper, we present DiffMSS, an object-focused conditional diffusion model designed to leverage semantic knowledge distillation for segmenting marine objects. Our model introduces a region-word matching mechanism to enable word-level selection of salient terms. These high-level textual semantic features are then utilized to guide the CFLN module in generating diffusive conditions through semantic knowledge distillation. To further enhance segmentation accuracy, we propose the CDS scheme, which effectively suppresses missegmentations of objects with fine-grained structures. Extensive experiments validate that DiffMSS surpasses the state-of-the-art methods in both quantitative and qualitative results.
Supplementary Material
In this supplementary material, we provide additional implementation details of the proposed DiffMSS, including its hyperparameter settings, training and testing procedures, precision-recall curves, discussion, and more visual comparison results.
6 Implementation Details
We detail the hyperparameter settings of DiffMSS in Table 8. The model is designed with an inner channel size of 64 and utilizes channel multipliers of [1, 2, 5, 8], enabling multi-scale feature extraction through hierarchical processing. The Adam optimizer is employed to adjust the learning rate, which is initially set to $1\times 10^{-4}$ and reduced by a factor of 0.8 every ten epochs. Each batch contains 32 samples with a resolution of $256\times 256$, balancing computational efficiency and spatial detail preservation.
[…]
To generate the text corresponding to each image, the pre-trained BLIP2 model [36] is used to initially produce descriptions of the image content. These descriptions are then calibrated by well-trained volunteers to correct inaccurate information. In the word-level semantic saliency extraction, the CLIP-ViT-B/32 model [58] is employed as both the image and text encoder, mapping multimodal features into a shared embedding space. The conditional feature learning network (CFLN) is initialized with PVTv2-B2 [69]. These carefully designed hyperparameters and methodologies enable DiffMSS to achieve accurate and efficient marine salient object segmentation.
Hyper-parameters Attributes
Conditional Channels 3
Inner Channels 64
Channel Multiplier [1, 2, 5, 8]
Number Workers 4
Batch Size 32
Patch Size 256 $\times$ 256
Epochs 150
Optimizer Adam
Learning Rate $1\times 10^{-4}$
Diffusion Steps 1000
Sampling Steps 10
Range of $\beta_{t}$ [0.0001,0.02]
Schedule of $\beta_{t}$ Linear
Clip Sample True
Pred Objective x0
Table 8: Hyper-parameters settings of the proposed DiffMSS.
7 Training and Testing
7.1 Training Procedure
The training inputs for DiffMSS consist of a degraded image $\mathcal{I}$, a corresponding caption $\mathcal{T}$ providing textual descriptions of the scene, and a reference saliency mask $\mathrm{x}_{0}$, which serves as the ground truth for supervising the segmentation task. These inputs collectively enable the model to effectively align visual and semantic features while learning to predict accurate segmentation masks.
Word-Level Semantic Saliency Extraction. The training process begins with word-level semantic saliency extraction, where salient terms describing objects are identified through word-region matching. This stage involves segmenting $M$ region masks $\mathbb{X}^{v}=\{\mathcal{X}_{i}^{v}\}_{i=1}^{M}$ and $N$ single-word tokens $\mathbb{X}^{t}=\{\mathcal{X}_{j}^{t}\}_{j=1}^{N}$. Visual and textual features, $\mathcal{F}_{i}^{v}$ and $\mathcal{F}_{j}^{t}$, are extracted from each region and word, respectively. The similarity between regions and words is calculated as $R_{k}=\frac{1}{MN}\sum_{i=1}^{M}\sum_{j=1}^{N}\text{Softmax}(\mathcal{F}_{i}^{v}{\mathcal{F}_{j}^{t}}^{T})$, which enables the identification of salient word tokens $\mathcal{F}^{st}$ that highlight key objects in the image.
Semantic Saliency Knowledge Distillation. The region-level semantic features $\mathcal{F}_{l}=\text{CFLN}(\mathcal{I},\mathbf{x}_{t},t)$ are derived and $\mathbf{x}_{t}$ represents the sample from the previous time step. These features act as conditions to the diffusion model and are optimized using a semantic-level consistency loss, $\mathcal{L}_{\text{consist}}=(\mathcal{F}_{l}^{v},\mathcal{F}^{st})$, to enhance the accurate representation of object saliency.
Object-Focused Diffusion and Sampling. The diffusion and sampling procedures employ conditional features to progressively generate object-focused saliency masks.
Algorithm 1 The training procedure of DiffMSS.
Require: Degraded image $\mathcal{I}$, corresponding caption $\mathcal{T}$, and reference saliency mask $\mathrm{x}_{0}$.
1:   repeat
2:      # Word-Level Semantic Saliency Extraction
3:      Segment $M$ region-level masks $\mathbb{X}^{v}=\left\{\mathcal{X}_{i}^{v}\right\}_{i=1}^{M}$
4:      Segment $N$ single-level words $\mathbb{X}^{t}=\left\{\mathcal{X}_{j}^{t}\right\}_{j=1}^{N}$
5:      Calculate the visual and textual features $\mathcal{F}_{i}^{v}=E_{v}(\mathcal{X}_{i}^{v})\mid_{i=1}^{M}$ and $\mathcal{F}_{j}^{t}=E_{t}(\mathcal{X}_{j}^{t})\mid_{j=1}^{N}$
6:      Calculate region-word similarity matching $R_{k}=\frac{1}{MN}\sum_{i=1}^{M}\sum_{j=1}^{N}Softmax\left(\mathcal{F}_{i}^{v}{\mathcal{F}_{j}^{t}}^{T}\right)$
7:      Identify saliency word-level tokens $\mathcal{F}^{st}=\left\{k\mid R\{k\}\geq m,\,k\in\left\{1,2,\dots,L_{t}\right\}\right\}$
8:      # Semantic Saliency Knowledge Distillation
9:      Calculate semantic conditional features $\mathcal{F}_{a}^{v}=CFLN(\mathcal{I},\mathbf{x}_{t},t)$
10:      Semantic-level knowledge constraints $\mathcal{L}_{\textit{consist}}=(\mathcal{F}_{l}^{v},\mathcal{F}^{st})$
11:      # Object-Focused Diffusion and Sampling
12:      $x_{0}\sim q(x_{0})$ and $\epsilon\sim\mathcal{N}(0,\textbf{I})$
13:      $t\sim\text{Uniform}\{1,2,\cdots,T\}$
14:      $\mathrm{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathrm{x}_{0}+(1-\bar{\alpha}_{t})\mathbf{I}$
15:      Sampling $\hat{\mathrm{x}}_{0}=\frac{\mathrm{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\,\epsilon_{\theta}(\mathrm{x}_{t},\mathcal{F}_{a}^{v},t)}{\sqrt{\bar{\alpha}_{t}}}$
16:      Perform gradient descent step on
$\nabla_{\theta}\left(\mathcal{L}_{consist}+\mathcal{L}_{\textit{BCE}}^{w}(\hat{\mathrm{x}}_{0},\mathrm{x}_{0})+\mathcal{L}_{\textit{IoU}}^{w}(\hat{\mathrm{x}}_{0},\mathrm{x}_{0})\right)$
17:   until converged
The model initializes with a sample $x_{0}\sim q(x_{0})$ and adds Gaussian noise $\epsilon\sim\mathcal{N}(0,\mathbf{I})$. At each time step $t$, the noisy sample $\mathbf{x}_{t}$ is progressively denoised as $\mathrm{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathrm{x}_{0}+(1-\bar{\alpha}_{t})\mathbf{I}$. The training process optimizes a joint loss function comprising the semantic consistency loss $\mathcal{L}_{\text{consist}}$, weighted binary cross-entropy loss $\mathcal{L}_{\text{BCE}}^{w}$, and weighted IoU loss $\mathcal{L}_{\text{IoU}}^{w}$, ensuring the accurate capture of salient objects. These steps constitute the forward diffusion process, and a comprehensive description of the training procedure is provided in Algorithm 1.
7.2 Testing Procedure
The testing procedure of DiffMSS requires a single degraded image $\mathcal{I}$ as input to generate segmentation results.
Algorithm 2 The testing procedure of DiffMSS.
Require: Degraded image $\mathcal{I}$.
1:   ${\mathrm{x}}_{T}\sim\mathcal{N}(0,\textbf{I})$
2:   ${P}_{t}=[\ ]$
3:   # Denoising Process
4:   for $t=T,\dots,1$ do
5:      Calculate semantic conditional features $\mathcal{F}_{a}^{v}=CFLN(\mathcal{I},\mathbf{x}_{t},t)$
6:      Sampling $\hat{\mathrm{x}}_{0}=\frac{\mathrm{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\,\epsilon_{\theta}(\mathrm{x}_{t},\mathcal{F}_{a}^{v},t)}{\sqrt{\bar{\alpha}_{t}}}$
7:      ${P}_{t}=Append(\hat{\mathbf{x}}_{0})$ Append $\hat{\mathbf{x}}_{0}$ to ${P}_{t}$
8:      $\mathbf{z}\sim\mathcal{N}(0,\mathbf{I})$ if $t>1$, else $\mathbf{z}=0$
9:      $\mathbf{x}_{t-1}=\frac{\sqrt{\alpha_{t}(1-\bar{\alpha}_{t-1})}}{1-\bar{\alpha}_{t}}\mathbf{x}_{t}+\frac{\sqrt{\bar{\alpha}_{t-1}\beta_{t}}}{1-\bar{\alpha}_{t}}\hat{\mathbf{x}}_{0}+\sigma_{t}\mathbf{z}$
10:   end for
11:   # Consensus Deterministic Sampling
12:   $M_{pre}=CDS({P}_{t})$
13:   return $M_{pre}$
The testing procedure begins by initializing a noise sample $\mathbf{x}_{T}\sim\mathcal{N}(0,\mathbf{I})$, which is iteratively refined through a denoising process. At each step $t$, conditional semantic features $\mathcal{F}_{a}^{v}=CFLN(\mathcal{I},\mathbf{x}_{t},t)$ are calculated and used to predict the clean reconstruction $\hat{\mathbf{x}}_{0}$. These predictions are stored in a history list ${\mathrm{P}}_{t}$, capturing the progressive refinements of the segmentation mask over iterations. For each time step $t$ from $T$ to 1, the predicted result $\hat{\mathbf{x}}_{0}$ is reconstructed using the model’s prediction $\epsilon_{\theta}(\mathbf{x}_{t},\mathcal{F}_{a}^{v},t)$, conditioned on the semantic features $\mathcal{F}_{a}^{v}$. The predicted mask $\hat{\mathbf{x}}_{0}$ is appended to ${\mathrm{P}}_{t}$. The next sample $\mathbf{x}_{t-1}$ is then computed by combining the noisy input $\mathbf{x}_{t}$, the predicted clean image $\hat{\mathbf{x}}_{0}$, and the linear noise schedule parameters, ensuring a gradual denoising process.
To further enhance the accuracy and robustness of the final segmentation, we employ the Consensus Deterministic Sampling (CDS) strategy that aggregates the intermediate predictions stored in ${\mathrm{P}}_{t}$ into a final output mask $M_{pre}$. This strategy improves the reliability and precision of the segmentation results without introducing additional computational overhead, as the intermediate predictions are naturally derived from the denoising process. The pseudo-code of this testing procedure is provided in Algorithm 2.
8 Analysis of Precision-Recall (PR) Curves
The Precision-Recall (PR) curve is a standard evaluation metric for segmentation models, which provides a comprehensive evaluation of a model’s ability to maintain accuracy (Precision) while capturing a broader range of relevant objects (Recall). Fig. 10 presents the Precision-Recall (PR) curves of various saliency segmentation models on the USOD10K, SUIM, and UFO-120 datasets. Across all datasets, our DiffMSS demonstrates superior performance, maintaining higher precision across a wide range of recall values compared to state-of-the-art methods such as DualSAM and RobustSAM.
Figure 10: Comparison of Precision-Recall (PR) curves for saliency segmentation models across three datasets: (a) USOD10K, (b) SUIM, and (c) UFO-120. These curves illustrate the superior performance of our DiffMSS through the Area-Under-the-Curve (AUC)-based analysis, which maintains high precision at most recall levels. Please zoom in to see more details.
Figure 11: More conditional features ablations of with (-w/ SKD) and without (-w/o SKD) Semantic Knowledge Distillation at different sampling steps $t$. The first row for each example shows the feature maps without SKD, while the second row presents those with SKD applied. The results highlight the enhanced feature representation and improved focus on salient regions achieved by SKD in various underwater scenes.
On all three datasets, DiffMSS achieves better segmentation accuracy, particularly at high recall levels whereas compared methods show a noticeable drop in precision. Specifically, for the USOD10K dataset, DiffMSS achieves the highest Area-Under-the-Curve (AUC), reflecting its superior segmentation accuracy in both high and moderate recall ranges. For the SUIM dataset, our method retains high precision even in the mid-recall range, outperforming RMFormer and BBRF. On the UFO-120 dataset, DiffMSS excels in capturing fine-grained object boundaries, maintaining precision where compared methods experience significant drops. The AUC-based analysis highlights the robustness of DiffMSS in capturing fine-grained details and effectively addressing the challenges of marine saliency segmentation tasks.
Figure 12: More visual comparison of segmentation results with (-w/ CDS) and without (-w/o CDS) Consensus Deterministic Sampling. The first column shows the original marine images, the second column presents the segmentation results without CDS, and the third column shows those with CDS applied. These results demonstrate that CDS can improve fine-grained segmentation.
9 More Visual Ablation Results
9.1 Visual Ablation of Conditional Features
To achieve object-focused conditional diffusion, we introduce Semantic Knowledge Distillation (SKD) to transfer the identified salient text features into the Conditional Feature Learning Network (CFLN) to generate conditions for object-focused diffusion. Fig. 11 illustrates a comparison of conditional feature maps with and without SKD at different sampling steps $t$. Each example consists of two rows: the first row shows feature maps without SKD, while the second row presents maps enhanced by SKD.
Figure 13: Our DiffMSS achieves more fine-grained visual segmentation results than Ground Truth (GT).
From the three examples, it is evident that SKD significantly sharpens feature representations and improves focus on salient regions. As the sampling step $t$ increases (from $t=0.9\mathrm{T}$ to $t=0$), the feature maps without SKD tend to lose focus and coherence. In contrast, the feature maps with SKD consistently maintain focus on salient regions, regardless of the sampling step. These results demonstrate that SKD not only improves feature clarity but also ensures robustness across different sampling stages, effectively addressing the challenges of underwater scenes, such as low contrast and cluttered backgrounds.
9.2 Visual Ablation of Consensus Deterministic Sampling
To enhance the segmentation of fine-grained anatomical structures in marine organisms, we introduce a Consensus Deterministic Sampling (CDS) strategy, which aggregates predictions from each denoising step. Fig. 12 presents a detailed comparison of segmentation results with (-w/ CDS) and without (-w/o CDS). The first column displays the raw underwater images, the second column shows segmentation results without CDS, and the third column highlights outcomes with CDS applied.
Figure 14: Visual comparison of our DiffMSS with the Ground Truth (GT). DiffMSS still segments several small objects that are not annotated in the GT, as marked by the yellow box.
Figure 15: Visual comparison of our DiffMSS and the Ground Truth (GT). Although DiffMSS achieves finer-grained segmentation of the main object compared to the GT, it loses some small camouflage objects, as marked by the yellow box.
The visual comparison demonstrates that CDS significantly improves fine-grained segmentation performance for marine instances. Specifically, without CDS (-w/o CDS), the segmentation results often fail to capture intricate object details, such as spines and thin boundaries, resulting in coarse or incomplete segmentation. In contrast, “-w/ CDS” effectively preserves these critical features and ensures accurate segmentation, particularly for marine objects with complex structures.
10 Limitation and Discussion
Although diffusion models with conditional prompts have demonstrated impressive performance in visual segmentation tasks, the potential to leverage contextual semantics for generating diffusion conditions remains underexplored. DiffMSS addresses this gap by extracting word-level semantic information from text descriptions and aligning it with visual representations. This alignment enables the diffusion model to capture fine-grained details and focus on key regions within the image.
Fine-grained Segmentation. Fig. 13 presents a visual comparison of DiffMSS and the Ground Truth (GT) masks. These results demonstrate that DiffMSS achieves finer-grained and more accurate segmentation compared to the reference. For objects with complex structures such as divers, corals, and lionfish, DiffMSS can effectively capture the intricate object boundaries and smaller details that are easily overlooked in manual annotations.
Fine-grained But Inaccurate Segmentation. We further evaluate the performance of DiffMSS in dealing with camouflaged small objects in complex underwater environments. As shown in Fig. 14, DiffMSS still segments several small objects that are not annotated in the GT. In addition, although DiffMSS achieves finer-grained segmentation of the main object compared to the GT, it loses some small camouflaged objects in Fig. 15. The reason for this issue is the inconsistency in how different volunteers handle small objects during manual annotation, leading to ambiguity in the dataset distribution. In other words, some underwater images include the segmentation of small objects, while others only segment the main objects and ignore the smaller ones. This variability in the dataset results in inaccurate segmentation outcomes for all segmentation methods, including but not limited to DiffMSS. In future work, we aim to propose a new dataset with standardized and consistent annotation criteria.
11 More Qualitative Comparison Results with State-of-the-Art Methods
To further demonstrate the superiority of DiffMSS over state-of-the-art salient object detection and segmentation methods, we present more qualitative results derived from publicly available benchmarks (USOD10K, SUIM, UFO-120). The comparison methods include SUIM-Net, RMFormer, BBRF, TC-USOD, RobustSAM, and DualSAM. Fig. 16 and Fig. 17 provide detailed visual comparisons of segmentation results for marine objects, including divers, shipwrecks, fish, and camouflaged marine species.
These comparison results encompass a variety of challenging underwater scenes, such as complex backgrounds and objects of varying sizes and appearances, highlighting the robustness of DiffMSS under diverse conditions. The results demonstrate that DiffMSS consistently outperforms compared methods, effectively capturing fine-grained object details and delivering superior segmentation performance across a wide range of marine saliencies.
Figure 16: More visualization results are presented by comparing our DiffMSS with state-of-the-art methods on the common underwater salient objects. The segmentation results are highlighted in orange for clarity.
Figure 17: More visualization results are presented by comparing our DiffMSS with state-of-the-art methods on the challenging underwater camouflage objects. The segmentation results are highlighted in orange for clarity.